Consider the following reaction: PbCO3(s)⇌PbO(s)+CO2(g)

Using data in Appendix C in the textbook, calculate the equilibrium pressure of CO2 in the system at 410 ∘C.

Using data in Appendix C in the textbook, calculate the equilibrium pressure of CO2 in the system at 220 ∘C.
Answer #1

Solution : Girven date : | 0 16 cos (S) 2 Pbo (s) + CO2 9) 1 Hoºc ; HO +213 = 683.5. AH = 88.3 +3/mot Pocos Pbo 8) + Cod (9)Poza é [-2.6482 re(2.6u8a) KPicoz. = 14.1285 @ T-980°c 220 +273 = H93 k. 4H = 8863 65/mol Pbcoz Pools) at coa 9 AH = 88.3 k388-3-74-5909 FP coa = e-( 88.B-74 46098802 - (13.7091 H.098803 sé 3.3446) tpicos 2.0.0352 Kenge 0.0352 196/